OpenStudy (anonymous):

(sinx/(1-cosx)) + (sinx/1+cosx)) = 2 cscx

OpenStudy (amoodarya):

\[\frac{sinx }{1-cosx}+\frac{sinx }{1+cosx}=\\sinx (\frac{1 }{1-cosx}+\frac{1 }{1+cosx})\\sinx (\frac{1+cosx +1-cosx }{1-\cos^2x})=\\ \frac {2\sin x}{1-\cos^2x}\] can you go on ?

OpenStudy (anonymous):

1-cos^2x=sin^2x 2sinx/sin^2x =2/sinx 2/sinx= 2csc x I think I get it thanks.